PHOTOCHEMISTRY OF MONOCHLORO COMPLEX OF COPPER (II) IN SOLUTION BY MEANS OF TRANSIENT ABSORPTION TIME-RESOLVED ULTRAFAST SPECTROSCOPY

\underline{A. S. MERESHCHENKO}, P. EL-KHOURY, S. PAL, A. N. TARNOVSKY, Bowling Green University, Department of Chemistry and Center for Photochemical Sciences, Bowling Green, Ohio 43403.

The photodissociation mechanism of copper (II) chloro complexes in methanol is studied by means of ultrafast broadband time-resolved deep-UV/UV/vis spectroscopy using [Cu^{II}(MeOH)_{5}Cl]^{+} as a model molecule. Upon 255 nm excitation, the majority of [Cu^{II}(MeOH)_{5}Cl]^{+} dissociate to form chlorine atoms Cl^{.} and penthacoordinate copper(I) solvated complexes [Cu^{I}(MeOH)_{5}]^{+}, which is followed by recombination back to the parent molecules. A three body-disscoation in which the nascent [Cu^{I}(MeOH)_{5}]^{+} fragment undergoes a methanol-ligand loss is observed to be of minor importance. The solvent influence on the photoreaction will be discussed.